OpenStudy (anonymous):

Help! should I use the law of cos than find the area? find the area of the triangle with the sides of 3,9, and 10

hero (hero):

Use Heron's formula: If you have a triangle with sides, a, b, c, then: \(A = \sqrt{s(s - a)(s - b)(s - c)}\) \(s = \dfrac{a + b + c}{2}\)

OpenStudy (anonymous):

Is there another way to do this?

hero (hero):

There's always more than one way to do any mathematics.

geerky42 (geerky42):

Well, what is another way to do this? @Hero

hero (hero):

 Find the height of \(d\), then you'd be able to use \(A_{\triangle} = \frac{bh}{2}\)